a pendulum which starts oscillation from the extreme position. A suppose that a bob is released from rest. determine the speed of the pendulum when passes position B and C. suppose that the length os the string OA is 1m and that the distance between point A and horizontal black line is 0.7m
1
Expert's answer
2019-05-29T09:30:02-0400

From the conservation of energy:


"mgh=0.5mv^2"

"v=\sqrt{2gh}"

"v=\sqrt{2(9.8)(0.7)}=3.7\frac{m}{s}"